# Practical solutions to the relative pose of three calibrated cameras

### v4

@@ 3.1 Approximate solutions to the 4p3v problem
In contrast to the (5,5,3) configuration, the (4,4,4) configuration, i.e., the configuration of four points in three views, leads to significantly more complex equations. State-of-the-art algebraic and numerical methods are known to fail in generating efficient and numerically stable solutions to these equations. Thus, solutions to the 4p3v problem require some approximations to be practical in real-world applications. This section introduces several practical approximate solutions to the 4p3v problem. Similarly to the 5pt+P3P solver, we decompose the problem into the problem of first estimating the relative pose of two cameras and then registering the third camera with an efficient P3P solver [41]. The idea is to assume an approximate geometry only in the first two views and use four input point correspondences to efficiently estimate this geometry. Next we describe two groups of such solvers.
@@
4p3v(A) solver: The first group of solvers approximates the geometry in the first two views using affine cameras. Although the approximation can be quite far from the correct geometry, [42] showed that an approximate affine camera model provides good accuracy for relative pose estimation if coupled with geometry refitting (using the full 7DoF fundamental matrix) and local optimization inside RANSAC. The proposed 4p3v(A) solver first uses four point correspondences in two views to efficiently estimate the affine fundamental matrix [18] and then registers the third camera using three triangulated points and a P3P solver [41].
@@
4p3v(M) solver: The affine camera model used in the 4p3v(A)-based solvers can be quite imprecise. This motivates us to introduce solvers that are based on a less restrictive approximation. In this case, we approximate only one point correspondence in two views. Under the assumption of a para-perspective projection, i.e., of affine geometry, it is known that the mean point of three 3D points is projected to the mean points of their projections in both images [59]. Thus, we generate one new approximate correspondence in two views as the correspondence between the mean points of three corresponding points detected in these views. Let be the mean point of three points , in the view , then is our new approximate correspondence. The new 4p3v(M) solver solves the 4p3v problem by first estimating the full 5DoF essential matrix from four original correspondences and the correspondence . This is done using the efficient 5pt relative pose solver [36]. As with previous solvers, the third camera is registered using a P3P solver [41]. In short, the 4p3v(M) solver solves the 4p3v problem using the 5pt+P3P solver and one approximate correspondence.
@@
As we show on large amounts of synthetic and real data (see Sec. 4 and Supp. mat. (SM)), the 4p3v(M) solver provides much more accurate estimates than the 4p3v(A) solver. Moreover, even in the basic form (without the modifications presented in the next section), on many scenes it performs comparably to the state-of-the-art HC solver [22]. This can be attributed to several facts and observations: (1) The correspondence does not need to be seen as a correspondence of points that are projections of the mean point of three 3D points and (in which case both and would have some error). We can look at this correspondence as a correspondence in which we fix a point in one view, e.g., , and generate a corresponding point in the second view. In this case, to generate a good correspondence, we only require the point in the second view to be reasonably close to the epipolar line defined by the mean point in the first view, i.e., the 2D point does not need to correspond to one particular 3D point with a given depth. (2) The ray from the center of the first camera through the mean point intersects the plane of and at point , which lies inside the triangle formed by these 3D points. The projection of into the second camera lies inside the triangle formed by . By construction, this projection lies on the epipolar line . Thus the epipolar line defined by passes through the triangle defined by . Consequently, the maximum distance of in the second image from the epipolar line is bounded by the maximum distance of from . For a formal lemma and proof, see SM. (3) For practical applications, when used in RANSAC, it is not necessary that each triplet generates a good correspondence . Samples with a high level of noise in the mean-point correspondence are filtered inside RANSAC.222This property was also used in the HC solver [22], which completely fails for many samples. These samples are filtered within RANSAC. On a large number of different scenes, we observed that even if some image pairs have triplets of points that generate very noisy mean-point correspondences, there are usually enough triplets for which the noise in is reasonably small to lead to good estimates. (4) Four point correspondences in two views usually fix the space of possible poses such that the correspondence, even if noisy, often generates a pose that is not very far from the ground truth pose. Such a pose is usually sufficient for filtering out outliers and a good initialization for non-linear optimization on the original four points in three views and subsequent local optimization on detected inliers. We support our observations by experiments on a large amount of data.
@@ 3.2 Making approximate solvers practical
The 4p3v(A)-based solvers, when used without any modifications, generally provide imprecise results even when used inside RANSAC with local optimization on three views. While the accuracy of the pure 4p3v(M) solver inside LO-RANSAC [31] (RANSAC with Local Optimization) is much better, there is still room for improvement. Here we present several simple modifications of these solvers that significantly boost their performance.
4p3v(M) solver: The mean point correspondence used in the 4p3v(M) solver can provide a good approximation of a correct correspondence. Yet, as mentioned above, it can also be noisy. In the 4p3v(M) solver, we thus, in addition to the mean point of three points in the second image, generate two additional points. These points are (1) if the longest dimension of the triangle is in the x-direction or (2) if it is in the y-direction. All three points, i.e., , , and are placed in correspondence with the mean point . The 4p3v(M) solver in the first step calls the 5pt solver [36] three times, with the correspondence being either , , or . The results of these three 5pt solvers are collected to create hypotheses for the relative pose of the first two cameras inside RANSAC. The shift is selected relative to the size of the triangle .
Early non-minimal refitting (ENM): The geometry estimated between the first two cameras using the proposed 4p3v(A) and 4p3v(M)- based solvers is only approximate. In standard LO-RANSAC, such a geometry is optimized in local optimization (LO) after registering the third view. However, this can lead to propagation of errors into triangulated points and subsequently into errors in the pose of the third camera. Inliers in three views w.r.t. such cameras together with imprecise pose initializations may not be sufficient for LO to converge to a good solution. We observed this especially for 4p3v(A)-based solvers. Fortunately, the design of our solvers allows us to optimize the geometry already after estimating the approximate geometry between the first two cameras. Here, even a very imprecise geometry returned by the 4p3v(A) solvers is usually sufficient to filter out outliers. Thus, after running the 4pt /5pt solver in the first step of the proposed solvers, we use the estimated approximate models to detect inliers in two views. We then refit the estimated geometry using the non-minimal version of the 5pt solver [36], which instead of a 4-dim null space of a matrix uses the last four vectors from the SVD/QR decomposition of a matrix.
@@ 4 Experiments
We extensively evaluated the proposed solvers on a large variety of synthetic and real data to test their robustness to noise, outliers, and scene properties, and to assess their performance inside state-of-the-art RANSAC-frameworks [2, 28]. We compare our novel solvers with the homotopy continuation 4p3v(HC) solver [22] and the 5pt+P3P baseline minimal solver for the (5,5,3) configuration.
Experimental setup. To obtain feature correspondences, we use SuperPoint [10] features with the LightGlue [33] matcher. We extract at most 2048 features per image. We perform matching for all three image pairs and keep only those matches that were consistently matched across all three views. We perform evaluation within two RANSAC frameworks: PoseLib [28] and GC-RANSAC [2]. For the 5pt solver, we use [36] and for the P3P solver, we use [41]. In PoseLib, we perform LO [7] using LM optimization. In GC-RANSAC, we perform LO using non-minimal solvers [36, 19] for fitting models to larger-than-minimal samples. We evaluated different shifts for our -based solvers using a validation scene (for the ablation study, see SM) and selected . The choice of is not critical, i.e., different values perform similarly.
Evaluation measures. Inspired by [23], we define the pose error as , where and are the angular errors of rotation and translation for pair in degrees [23]. We also report AUC values [23] at different thresholds for the pose error. We include results for an alternative pose error definition which includes and in SM.
@@
An approximation error introduced by the para-perspective projection, i.e., the affine geometry, and its effect on absolute camera pose determination is studied in the literature [59, 21]. Here, we thus only study the effect of the proposed approximation on relative pose estimation for real-world data and interesting synthetic scenarios. In the following experiments, all solvers are tested on two views without RANSAC. The two-view variants of our solvers are denoted as 4p(A), 4p(M), and 4p(M).
@@
Similar observations as on synthetic data can be made on real-world data. Tab. 1 shows results on different scenes from the PhotoTourism dataset [23]. As can be seen, the approximate solvers are not as accurate as the 5pt solver when used without ENM. The gap between the 4p(M)/4p(M) solvers and the 5pt solver is for some scenes larger (Sacre Coeur, Trevi Fountain), and for some noticeably smaller (Reichstag, St. Peter’s Square), showing that the performance of our solvers is scene-dependent. The 4p(A) solver returns large errors. The accuracy of the 4p(M)/4p(M) solvers is not very far from the 5pt solver, with the 4p(M) solver even slightly outperforming the 5pt solver on some scenes (Reichstag, Taj Mahal). ENM helps to increase the precision of all solvers, especially 4p(A). Still, the 4p(A)+ENM solver returns quite high errors. However, in practice, all solvers are used inside RANSAC, where the current best pose is locally optimized on inliers. LO compensates for less accurate pose estimates and, as we show on real experiments for three as well as two views (see SM), it can suppress the errors of the 4p(A)/4p3v(A)-based solvers, making them comparable to the 4p(M)/4p3v(M)-based solvers. For the three view scenario, the performance of all of our solvers can be further improved by pose refinement (+R) and filtering (+F).
@@
In the last set of experiments, we study the accuracy of the mean point correspondence. We sample 100 four-tuples of point correspondences consistent with the ground truth relative pose, i.e., inliers, for each image pair in scenes from the PhotoTourism dataset [23]. We use the first three correspondences to define the triangles in both images. Then, we establish correspondences between the mean of the triangle in one image and various points in the triangle in the second image. We express points in the second triangle via their barycentric coordinates and uniformly sample barycentric coordinates , such that .
@@
Experiments on real data. We test the solvers on all scenes from the PhotoTourism dataset [52, 23] which provide ground truth poses and intrinsics via a COLMAP [50] reconstruction. In the results, we do not include the St. Peter’s Square scene that we used for the validation of and the number of refinement iterations (see SM). We also include results for the Cambridge Landmarks dataset [24] (except the Street scene, which is commonly not used due to issues with its ground truth) and Aachen Day-Night v1.1 [60]. For PhotoTourism and Aachen, we use the images in their original resolution. For Cambridge Landmarks, we resize them so that the larger side is 800 px. For each scene, we sample 5,000 random image triplets with at least 10 matches obtained with [10, 33] with at least overlap [23].
| PhotoTourism [23] | Cambridge Landmarks [24] | Aachen Day-Night v1.1 [60] | ||||||||||||||||
@@
Tab. 2 shows the results for PoseLib RANSAC with early termination and 5px epipolar threshold. Results for GC-RANSAC are in SM. With the suggested modifications (see Sec. 3.2) all proposed solvers outperform the state-of-the-art 4p3v(HC) solver [22] in terms of pose accuracy with many of the variants also achieving faster runtimes.
@@
We also investigate the speed-accuracy trade-off of the solvers by running PoseLib RANSAC [28] for a set of fixed numbers of iterations. Runtimes are reported for 1 core of a 2 GHz Intel Xeon Gold 6338 CPU. As shown in Fig. 4, the proposed 4p3v(M)+R+F solver consistently provides the best speed-accuracy trade-off both with and without ENM. 4p3v(A)+R+F+ENM provides a similar performance, typically beating 4p3v(HC) [22]. However, as shown in Fig. 4(d) for the St. Mary’s Church scene [24], it may perform worse for some specific scenes.
@@
Limitations. As discussed, the accuracy of the proposed approximate solvers is scene dependent.444This weakness also applies to [22], since the scene needs to be similar enough to the training scenes for the MLP-based classifier to work well. However, after using the propsed ENM refitting, the scene-dependency is negligible. Especially for the 4p3v(M)+ENM solver, we have not noticed performance drops for some specific scene geometries or camera configurations. While the two-view variants of our solvers 4p(M)/4p(M) provide very similar results to the 5pt solver in two views, in this case they do not outperform the 5pt solver. In the three-view scenario, the better performance is achieved thanks to the proposed pose refinement (+R) and filtering (+F).
The proposed ENM refitting cannot be applied to the 4p3v(HC) solver [22], since this solver estimates the pose of all three cameras together. It is theoretically applicable to the solver proposed in [37]. However, here ENM would have needed to be run on hundreds of candidate poses corresponding to tens-to-hundreds of sampled epipoles. This is because for one fixed point on the curve of epipoles, the error of the sampled epipole is not bounded since the true epipole can lie anywhere on the curve, even outside the image. Thus, ENM might avoid local minima present in the second step of the original solver, but would significantly slow down the solver in the first step, making it much slower than our 4p3v(M)+ENM/ 4p3v(M)+ENM solvers.

### v5

@@ 3.1 Approximate solutions to the 4p3v problem
In contrast to the (5,5,3) configuration, the (4,4,4) configuration, i.e., the configuration of four points in three views, leads to significantly more complex equations. State-of-the-art algebraic and numerical methods are known to fail in generating efficient and numerically stable solutions to these equations. Thus, solutions to the 4p3v problem require some approximations to be practical in real-world applications. This section introduces several practical approximate solutions to the 4p3v problem. Similarly to the 5pt+P3P solver, we decompose the problem into the problem of first estimating the relative pose of two cameras and then registering the third camera with an efficient P3P solver [lambda-twist]. The idea is to assume an approximate geometry only in the first two views and use four input point correspondences to efficiently estimate this geometry. Next we describe two groups of such solvers.
@@
4p3v(A) solver: The first group of solvers approximates the geometry in the first two views using affine cameras. Although the approximation can be quite far from the correct geometry, [pritts_ivcnz13] showed that an approximate affine camera model provides good accuracy for relative pose estimation if coupled with geometry refitting (using the full 7DoF fundamental matrix) and local optimization inside RANSAC. The proposed 4p3v(A) solver first uses four point correspondences in two views to efficiently estimate the affine fundamental matrix [hartley2006multiple] and then registers the third camera using three triangulated points and a P3P solver [lambda-twist].
@@
4p3v(M) solver: The affine camera model used in the 4p3v(A)-based solvers can be quite imprecise. This motivates us to introduce solvers that are based on a less restrictive approximation. In this case, we approximate only one point correspondence in two views. Under the assumption of a para-perspective projection, i.e., of affine geometry, it is known that the mean point of three 3D points is projected to the mean points of their projections in both images [Zhang2014]. Thus, we generate one new approximate correspondence in two views as the correspondence between the mean points of three corresponding points detected in these views. Let be the mean point of three points , in the view , then is our new approximate correspondence. The new 4p3v(M) solver solves the 4p3v problem by first estimating the full 5DoF essential matrix from four original correspondences and the correspondence . This is done using the efficient 5pt relative pose solver [Nister-5pt-PAMI-2004]. As with previous solvers, the third camera is registered using a P3P solver [lambda-twist]. In short, the 4p3v(M) solver solves the 4p3v problem using the 5pt+P3P solver and one approximate correspondence.
@@
As we show on large amounts of synthetic and real data (see Sec. 4 and Supp. mat. (SM)), the 4p3v(M) solver provides much more accurate estimates than the 4p3v(A) solver. Moreover, even in the basic form (without the modifications presented in the next section), on many scenes it performs comparably to the state-of-the-art HC solver [Hruby_cvpr2022]. This can be attributed to several facts and observations: (1) The correspondence does not need to be seen as a correspondence of points that are projections of the mean point of three 3D points and (in which case both and would have some error). We can look at this correspondence as a correspondence in which we fix a point in one view, e.g., , and generate a corresponding point in the second view. In this case, to generate a good correspondence, we only require the point in the second view to be reasonably close to the epipolar line defined by the mean point in the first view, i.e., the 2D point does not need to correspond to one particular 3D point with a given depth. (2) The ray from the center of the first camera through the mean point intersects the plane of and at point , which lies inside the triangle formed by these 3D points. The projection of into the second camera lies inside the triangle formed by . By construction, this projection lies on the epipolar line . Thus the epipolar line defined by passes through the triangle defined by . Consequently, the maximum distance of in the second image from the epipolar line is bounded by the maximum distance of from . For a formal lemma and proof, see SM. (3) For practical applications, when used in RANSAC, it is not necessary that each triplet generates a good correspondence . Samples with a high level of noise in the mean-point correspondence are filtered inside RANSAC.222This property was also used in the HC solver [Hruby_cvpr2022], which completely fails for many samples. These samples are filtered within RANSAC. On a large number of different scenes, we observed that even if some image pairs have triplets of points that generate very noisy mean-point correspondences, there are usually enough triplets for which the noise in is reasonably small to lead to good estimates. (4) Four point correspondences in two views usually fix the space of possible poses such that the correspondence, even if noisy, often generates a pose that is not very far from the ground truth pose. Such a pose is usually sufficient for filtering out outliers and a good initialization for non-linear optimization on the original four points in three views and subsequent local optimization on detected inliers. We support our observations by experiments on a large amount of data.
@@ 3.2 Making approximate solvers practical
The 4p3v(A)-based solvers, when used without any modifications, generally provide imprecise results even when used inside RANSAC with local optimization on three views. While the accuracy of the pure 4p3v(M) solver inside LO-RANSAC [lebedaLO] (RANSAC with Local Optimization) is much better, there is still room for improvement. Here we present several simple modifications of these solvers that significantly boost their performance.
4p3v(M) solver: The mean point correspondence used in the 4p3v(M) solver can provide a good approximation of a correct correspondence. Yet, as mentioned above, it can also be noisy. In the 4p3v(M) solver, we thus, in addition to the mean point of three points in the second image, generate two additional points. These points are (1) if the longest dimension of the triangle is in the x-direction or (2) if it is in the y-direction. All three points, i.e., , , and are placed in correspondence with the mean point . The 4p3v(M) solver in the first step calls the 5pt solver [Nister-5pt-PAMI-2004] three times, with the correspondence being either , , or . The results of these three 5pt solvers are collected to create hypotheses for the relative pose of the first two cameras inside RANSAC. The shift is selected relative to the size of the triangle .
Early non-minimal refitting (ENM): The geometry estimated between the first two cameras using the proposed 4p3v(A) and 4p3v(M)- based solvers is only approximate. In standard LO-RANSAC, such a geometry is optimized in local optimization (LO) after registering the third view. However, this can lead to propagation of errors into triangulated points and subsequently into errors in the pose of the third camera. Inliers in three views w.r.t. such cameras together with imprecise pose initializations may not be sufficient for LO to converge to a good solution. We observed this especially for 4p3v(A)-based solvers. Fortunately, the design of our solvers allows us to optimize the geometry already after estimating the approximate geometry between the first two cameras. Here, even a very imprecise geometry returned by the 4p3v(A) solvers is usually sufficient to filter out outliers. Thus, after running the 4pt /5pt solver in the first step of the proposed solvers, we use the estimated approximate models to detect inliers in two views. We then refit the estimated geometry using the non-minimal version of the 5pt solver [Nister-5pt-PAMI-2004], which instead of a 4-dim null space of a matrix uses the last four vectors from the SVD/QR decomposition of a matrix.
@@ 4 Experiments
We extensively evaluated the proposed solvers on a large variety of synthetic and real data to test their robustness to noise, outliers, and scene properties, and to assess their performance inside state-of-the-art RANSAC-frameworks [barath2017graph, PoseLib]. We compare our novel solvers with the homotopy continuation 4p3v(HC) solver [Hruby_cvpr2022] and the 5pt+P3P baseline minimal solver for the (5,5,3) configuration.
Experimental setup. To obtain feature correspondences, we use SuperPoint [detone2018superpoint] features with the LightGlue [lindenberger2023lightglue] matcher. We extract at most 2048 features per image. We perform matching for all three image pairs and keep only those matches that were consistently matched across all three views. We perform evaluation within two RANSAC frameworks: PoseLib [PoseLib] and GC-RANSAC [barath2017graph]. For the 5pt solver, we use [Nister-5pt-PAMI-2004] and for the P3P solver, we use [lambda-twist]. In PoseLib, we perform LO [Chum-2003] using LM optimization. In GC-RANSAC, we perform LO using non-minimal solvers [Nister-5pt-PAMI-2004, dlspnp] for fitting models to larger-than-minimal samples. We evaluated different shifts for our -based solvers using a validation scene (for the ablation study, see SM) and selected . The choice of is not critical, i.e., different values perform similarly.
Evaluation measures. Inspired by [IMC2020], we define the pose error as , where and are the angular errors of rotation and translation for pair in degrees [IMC2020]. We also report AUC values [IMC2020] at different thresholds for the pose error. We include results for an alternative pose error definition which includes and in SM.
@@
An approximation error introduced by the para-perspective projection, i.e., the affine geometry, and its effect on absolute camera pose determination is studied in the literature [Zhang2014, Horaud97]. Here, we thus only study the effect of the proposed approximation on relative pose estimation for real-world data and interesting synthetic scenarios. In the following experiments, all solvers are tested on two views without RANSAC. The two-view variants of our solvers are denoted as 4p(A), 4p(M), and 4p(M).
@@
Similar observations as on synthetic data can be made on real-world data. Tab. 1 shows results on different scenes from the PhotoTourism dataset [IMC2020]. As can be seen, the approximate solvers are not as accurate as the 5pt solver when used without ENM. The gap between the 4p(M)/4p(M) solvers and the 5pt solver is for some scenes larger (Sacre Coeur, Trevi Fountain), and for some noticeably smaller (Reichstag, St. Peter’s Square), showing that the performance of our solvers is scene-dependent. The 4p(A) solver returns large errors. The accuracy of the 4p(M)/4p(M) solvers is not very far from the 5pt solver, with the 4p(M) solver even slightly outperforming the 5pt solver on some scenes (Reichstag, Taj Mahal). ENM helps to increase the precision of all solvers, especially 4p(A). Still, the 4p(A)+ENM solver returns quite high errors. However, in practice, all solvers are used inside RANSAC, where the current best pose is locally optimized on inliers. LO compensates for less accurate pose estimates and, as we show on real experiments for three as well as two views (see SM), it can suppress the errors of the 4p(A)/4p3v(A)-based solvers, making them comparable to the 4p(M)/4p3v(M)-based solvers. For the three view scenario, the performance of all of our solvers can be further improved by pose refinement (+R) and filtering (+F).
@@
In the last set of experiments, we study the accuracy of the mean point correspondence. We sample 100 four-tuples of point correspondences consistent with the ground truth relative pose, i.e., inliers, for each image pair in scenes from the PhotoTourism dataset [IMC2020]. We use the first three correspondences to define the triangles in both images. Then, we establish correspondences between the mean of the triangle in one image and various points in the triangle in the second image. We express points in the second triangle via their barycentric coordinates and uniformly sample barycentric coordinates , such that .
@@
Experiments on real data. We test the solvers on all scenes from the PhotoTourism dataset [snavely2006photo, IMC2020] which provide ground truth poses and intrinsics via a COLMAP [Schoenberger2016CVPR] reconstruction. In the results, we do not include the St. Peter’s Square scene that we used for the validation of and the number of refinement iterations (see SM). We also include results for the Cambridge Landmarks dataset [kendall2015cambridge] (except the Street scene, which is commonly not used due to issues with its ground truth) and Aachen Day-Night v1.1 [zhang2021aachen]. For PhotoTourism and Aachen, we use the images in their original resolution. For Cambridge Landmarks, we resize them so that the larger side is 800 px. For each scene, we sample 5,000 random image triplets with at least 10 matches obtained with [detone2018superpoint, lindenberger2023lightglue] with at least overlap [IMC2020].
| PhotoTourism [IMC2020] | Cambridge Landmarks [kendall2015cambridge] | Aachen Day-Night v1.1 [zhang2021aachen] | ||||||||||||||||
@@
| PhotoTourism [IMC2020] | Cambridge Landmarks [kendall2015cambridge] | Aachen Day-Night v1.1 [zhang2021aachen] | ||||||||||||||||
| 4p3v(HC) [Hruby_cvpr2022] | 6.37 | 1.62 | 62.41 | 73.53 | 81.94 | 64.10 | 8.73 | 2.67 | 47.99 | 64.58 | 76.65 | 60.11 | 12.43 | 3.55 | 42.28 | 55.20 | 66.55 | 67.26 |
[…]
Tab. 3 shows the results for PoseLib RANSAC with early termination and 5px epipolar threshold. Results for GC-RANSAC are in SM. With the suggested modifications (see Sec. 3.2) all proposed solvers outperform the state-of-the-art 4p3v(HC) solver [Hruby_cvpr2022] in terms of pose accuracy with many of the variants also achieving faster runtimes.
@@
We also investigate the speed-accuracy trade-off of the solvers by running PoseLib RANSAC [PoseLib] for a set of fixed numbers of iterations. Runtimes are reported for 1 core of a 2 GHz Intel Xeon Gold 6338 CPU. As shown in Fig. 4, the proposed 4p3v(M)+R+F solver consistently provides the best speed-accuracy trade-off both with and without ENM. 4p3v(A)+R+F+ENM provides a similar performance, typically beating 4p3v(HC) [Hruby_cvpr2022]. However, as shown in Fig. 4(d) for the St. Mary’s Church scene [kendall2015cambridge], it may perform worse for some specific scenes.
@@
Limitations. As discussed, the accuracy of the proposed approximate solvers is scene dependent.444This weakness also applies to [Hruby_cvpr2022], since the scene needs to be similar enough to the training scenes for the MLP-based classifier to work well. However, after using the propsed ENM refitting, the scene-dependency is negligible. Especially for the 4p3v(M)+ENM solver, we have not noticed performance drops for some specific scene geometries or camera configurations. While the two-view variants of our solvers 4p(M)/4p(M) provide very similar results to the 5pt solver in two views, in this case they do not outperform the 5pt solver. In the three-view scenario, the better performance is achieved thanks to the proposed pose refinement (+R) and filtering (+F).
The proposed ENM refitting cannot be applied to the 4p3v(HC) solver [Hruby_cvpr2022], since this solver estimates the pose of all three cameras together. It is theoretically applicable to the solver proposed in [DBLP:journals/ijcv/NisterS06]. However, here ENM would have needed to be run on hundreds of candidate poses corresponding to tens-to-hundreds of sampled epipoles. This is because for one fixed point on the curve of epipoles, the error of the sampled epipole is not bounded since the true epipole can lie anywhere on the curve, even outside the image. Thus, ENM might avoid local minima present in the second step of the original solver, but would significantly slow down the solver in the first step, making it much slower than our 4p3v(M)+ENM/ 4p3v(M)+ENM solvers.